# Union of Finitely Generated Congruences on Ground Term Algebra

### v1

@@ 16 Main Result
Let $E$ and $F$ be GTESs over a signature $\Sigma$. It is decidable in $O(n^{2})$ time whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$, where $n=size(E)+size(F)$.
@@
  • $BSTEP\langle E\cup F,\emptyset\rangle$ stored in a red-black tree saved in a variable $RBT6$; for every ordered pair $(a,b)\in BSTEP\langle E\cup F,\emptyset\rangle$, the search key is the first component $a$, the second component, $b$ is satellite data;
@@
  • the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$, and for every vertex $a\in C\langle E\cup F,\emptyset\rangle$, the values of the attributes $a.equal_{E}$, $a.equal_{F}$, $a.keeps_{E}$, and $a.keeps_{F}$.
@@
Consequence 16.140.
Let $E$ and $F$ be GTESs over a signature $\Sigma$. Then we can decide in $O(n^{2})$ time whether there exists a GTES $H$ over $\Sigma$ such that $\mathop{\leftrightarrow}^{*}_{H}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$, where $n=size(E)+size(F)$. If the answer is ’yes’, then $\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ is a congruence on the ground term algebra ${\cal T}(\Sigma)$ and $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. Otherwise $\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ is not a congruence on the ground term algebra ${\cal T}(\Sigma)$.
@@
By Theorem 16.139, we decide in $O(n^{2})$ time whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. By Proposition 9.82, if the answer is ’yes’, then there exists a GTES $H$ over $\Sigma$ such that $\mathop{\leftrightarrow}^{*}_{H}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$, and $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$; otherwise $\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ is not a congruence on the ground term algebra ${\cal T}(\Sigma)$. ∎
@@ 17 Conclusion
We showed that we can decide in $O(n^{2})$ time for given GTESs $E$ and $F$ over a signature $\Sigma$ whether Statement 1 above holds. Here $n=size(E)+size(F)$. We can construct the GTES $E\cup F$ in $O(n^{2})$ time.
@@
Given GTES $E$ and GTES $F$, for every $p\in W$, we can construct the deterministic bottom-up tree automata ${\cal A}$, ${\cal B}$, and ${\cal C}$ such that ${\cal A}$ recognizes $p/_{\mathop{\leftrightarrow}^{*}_{E}}$, ${\cal B}$ recognizes $p/_{\mathop{\leftrightarrow}^{*}_{F}}$, and ${\cal C}$ recognizes $p/_{\mathop{\leftrightarrow}^{*}_{E}\cup F}$, see Proposition 2.10. Thence, in Main Cases 1, 2, and 3, we can reduce our decidability problem to deciding the equivalence of two deterministic bottom-up tree automata. Champavère et al [5] presented an algorithm for deciding inclusion between two tree languages recognized by given bottom-up tree automata. Applying their algorithm to our decision problem in Main Cases 1, 2, and 3 yields an alternative decision algorithm, which is more sophisticated than ours, and in general takes more time than ours, see Lemmas 12.103, 13.116, and 14.121.
@@
Consider the unary ranked alphabet $\Sigma=\Sigma_{0}=\{\,\#,\;\$,\;\mathsterling\,\}$. Let
$D=\{\,\#\doteq\$\,\}$,  $E=\{\,\#\doteq\mathsterling\,\}$,  $F=\{\,\$\doteq\mathsterling\,\}$,  $H=D\cup E\cup F$.
@@
  • #$\#/_{\mathop{\leftrightarrow}^{*}_{D}}\subset\#/_{\mathop{\leftrightarrow}^{*}_{H}}$,   #$\#/_{\mathop{\leftrightarrow}^{*}_{E}}\subset\#/_{\mathop{\leftrightarrow}^{*}_{H}}$,   #$\#/_{\mathop{\leftrightarrow}^{*}_{F}}\subset\#/_{\mathop{\leftrightarrow}^{*}_{H}}$,
  • #$\#/_{\mathop{\leftrightarrow}^{*}_{D}},\;\#/_{\mathop{\leftrightarrow}^{*}_{E}},\;\#/_{\mathop{\leftrightarrow}^{*}_{F}}$ are pairwise incomparable for the inclusion relation, and

### v2

@@ 16 Main Result
Let $E$ and $F$ be GTESs over a signature $\Sigma$. It is decidable in $O(n^{2})$ time whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$, where $n=\mathrm{size}(E)+\mathrm{size}(F)$.
@@
  • 8.
    $BSTEP\langle E\cup F,\emptyset\rangle$ stored in a red-black tree saved in a variable $RBT6$; for every ordered pair $({b},{c})\in BSTEP\langle E\cup F,\emptyset\rangle$, the search key is the first component b{b}, the second component, c{c} is satellite data;
  • 9.
@@
  • 11.
    the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$, and for every vertex ${b}\in C\langle E\cup F,\emptyset\rangle$, the values of the attributes ${b}.equal_{E}$, ${b}.equal_{F}$, ${b}.keeps_{E}$, and ${b}.keeps_{F}$.
@@
Consequence 16.150.
Let $E$ and $F$ be GTESs over a signature $\Sigma$. Then we can decide in $O(n^{2})$ time whether there exists a GTES $H$ over $\Sigma$ such that $\mathop{\leftrightarrow}^{*}_{H}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$, where $n=\mathrm{size}(E)+\mathrm{size}(F)$. If the answer is “yes”, then $\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ is a congruence on the ground term algebra ${\cal T}(\Sigma)$ and $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. Otherwise $\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ is not a congruence on the ground term algebra ${\cal T}(\Sigma)$.
@@
By Theorem 16.149, we decide in $O(n^{2})$ time whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. By Proposition 9.92, if the answer is “yes”, then there exists a GTES $H$ over $\Sigma$ such that $\mathop{\leftrightarrow}^{*}_{H}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$, and $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$; otherwise $\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ is not a congruence on the ground term algebra ${\cal T}(\Sigma)$. ∎
@@ 17 Conclusion
We showed that we can decide in $O(n^{2})$ time for given GTESs $E$ and $F$ over a signature $\Sigma$ whether Statement 1 above holds. Here $n=\mathrm{size}(E)+\mathrm{size}(F)$. We can construct the GTES $E\cup F$ in $O(n^{2})$ time.
@@
[…]
@@
[…]
$D=\{\,\#\doteq\mathdollar\,\}$,  $E=\{\,\#\doteq\mathsterling\,\}$,  $F=\{\,\mathdollar\doteq\mathsterling\,\}$,  $H=D\cup E\cup F$.
@@
  • 3.
    $_{\mathop{\leftrightarrow}^{*}_{D}}\subset\#/_{\mathop{\leftrightarrow}^{*}_{H}}$,   $_{\mathop{\leftrightarrow}^{*}_{E}}\subset\#/_{\mathop{\leftrightarrow}^{*}_{H}}$,   $_{\mathop{\leftrightarrow}^{*}_{F}}\subset\#/_{\mathop{\leftrightarrow}^{*}_{H}}$,
  • 4.
    $_{\mathop{\leftrightarrow}^{*}_{D}},\;\#/_{\mathop{\leftrightarrow}^{*}_{E}},\;\#/_{\mathop{\leftrightarrow}^{*}_{F}}$ are pairwise incomparable for the inclusion relation, and
  • 5.